# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ IV.5 Pulse length dependence
In most applications, the electromagnetic field driving the breakdown is a pulsed RF signal. Thus, the driving forces act for a limited pulse length, , which allows for exploration of the kinetics of the breakdown nucleation driving force. Assuming that the plastic activity within the sample follows stochastic evolution leads to a strong dependence of the BDR on , which can be empirically shown to satisfy
@@
see Fig. 44 [45]. Here, is the BDR, and , , , and are constants depending on the field. Such behavior was measured in connection with the CLIC project where pulse lengths between 50 and 400 ns were examined, with a duty cycle of 50 Hz.[184] Measurements of the dependence of breakdown rate on pulse length are described in Sec. III.5.
@@
In various scenarios, the BDR was shown experimentally, to have an exponential or power-law dependence on .[39, 58, 31] However, the validity of using the existing data to determine the dependence is limited, as it consists of either a small sample,[39] or of measurements taken during the conditioning process, when the BDR is still dominated by extrinsic processes.[31] At the very least, the BDR is expected to saturate for a continuous-wave RF signal, and therefore, the exponential or power-law dependence holds only for a limited range of pulse lengths.
@@
Under the assumption of inter-pulse independence, as discussed in III with relevant data shown in Fig. 16, the distribution of breakdowns in time within each pulse should be an increasing function, due to the finite evolution of the time to transition . Measurements of the distribution of time of breakdowns within a pulse are described in Sec. III.3. Figure 45 shows this distribution for = 250 MV/m and a pulse duration of 400 ns. For a given time interval within a pulse, an event will mature if it began within the time interval . Given that a breakdown occurred, the probability that it occurred within an interval is, therefore, , where is linearly proportional to the probability distribution function. This integral, however, is simply the CDF of , which is derived from the MDDF model. This non-Poissonian distribution becomes predominantly Poissonian for times that are significantly greater than .
@@
If, however, the interval between pulses is smaller than the typical relaxation time, then the breakdown probability should not depend on the pulse duration alone, but rather on the combined effect of exposure to the field and the relaxation achieved between pulses. In this case, the variation in the BDR within the pulse can be small, and characterized by a constant probability, similarly to the slow variation observed for ns in Fig. 45. Indeed, in Ref. [185] it was shown that the breakdown distribution does not vary significantly within the pulse. However, an increase in the breakdown probability was observed for one of the structures in Ref. [185], and it was not clear whether full conditioning was achieved. Therefore, experiments involving variation of the duty cycle, as well as further data regarding the pulse-length dependence during and after conditioning, can help determine the exact nature of the pulse length dependence of the BDR and help determine the memory effect between and within pulses.
@@
The success of dislocation-based models in describing distributions of BD, and pre-BD phenomena together with the consistent microscopic observations provide significant support to the hypothesis that indeed dislocation-based dynamics control the early stages prior to plasma onset.
@@ IV.6 Other indications of plastic activity
Significant dislocation activity can lead to various signals, including electric signals, due to effects on resistivity and acoustic emission due to collision of dislocations with obstacles. In these cases, macroscopic signals help quantify variations in the global density of dislocations or significant increases in the number of moving dislocations. such measurement independent of microscopic observations may help corroborate the assumption that dislocation motion is related to the observed processes. For plastically induced cyclic loading it was shown previously that acoustic emission by moving dislocations produced a measurable signal [97] However efforts to measure the acoustic emission for surfaces exposed to electric fields have not produced until now measurable signals. Additional Examples of such processes are effects of the dislocations density on local variation of resistivity in the top layer which is expected to affect the resistivity of the uppermost grains and is expected to yield a measurable signal such as was measured in direct plastic deformation [117]. For both methods, initial trials were performed, and for the resistivity system detailed design was simulated [26], but as of now the mensurability of pre-BD signals remains an unsolved issue.
@@ V Surface dynamics
Surfaces subject to high electric fields undergo several dynamic processes that can create and/or amplify various surface irregularities. As it was shown in Chapter IV, some few atomic layer protrusions, step edges, or even larger irregularities may develop on the surface as a result of under-the-surface dislocation dynamics triggered by the Maxwell stress that is a consequence of the application of an external electric field.
@@
Strictly speaking, the interatomic forces that hold the material together at the surface act towards smoothening any surface irregularity following the principle of surface energy minimization [22]. This implies that any atomically sharp surface feature experiences a strong thermodynamic force towards a blunter and, hence, energetically more stable shape. This effect was observed in kinetic Monte Carlo atomistic simulations in Ref. [65]. However, the features resulting from dislocation activity under the surface can develop further due to forces of electrostatic nature acting on the polarized and partially charged surface atoms in the presence of an external electric field.
@@
There is significant experimental evidence that in the presence of an external electric field the diffusion of surface atoms becomes biased towards the field gradients (i.e. inhomogeneity of the electric field around surface irregularities) [170, 169, 73]. This effect was shown to lead to the formation of surface features, e.g. the nanoprotrusions were seen to grow in several experimental studies [16, 122, 54, 187, 109].
@@
Fig. 46a illustrates the faceted surface of a W tip exposed for 5 hours to the simultaneous effect of an electric field and local heating by a laser [187]. The inset shows the field emission map after 2.5 hours since the beginning of the experiment. Here the bright spots correspond to the localized field emission currents due to the appearance of field-emitting spots along the edges of the surface facets. Weak emission is still visible from the ridges, which was stronger at the beginning of the experiment but started fading away when more localized field-emitting spots started to grow. Eventually, several initial spots merged into a single one at the sharpest corner (similar to that shown in Fig. 46b) with a much stronger current at a lower applied voltage. We note that the W tip was carefully relaxed before the experiment to ensure the absence of any irregularity on the tip surface, which could trigger the development of the observed faceted structures with the subsequent formation of field-emitting spots. The significant reduction of the applied voltage (up to three times) observed in this experiment, which still led to a strong field emission signal cannot be explained by any other phenomenon. The corresponding finite-element method (FEM) calculations, see Fig. 46b, confirmed that similar field enhancement (a.k.a. the applied voltage reduction) can be obtained if one assumes the formation of a protrusion at the corner of a surface facet. The latter may form as a result of thermal stresses during the rapid heating and cooling of the surface by femtosecond laser irradiation. Here, the color bar indicates the electrostatic field value with the highest value changing from left to right from 2 V/nm to 7.34 V/nm, measured at the small corner protrusion.
@@ V.1 Theoretical modeling of surface charge effects on atomic dynamics
From the Electrodynamics point of view, metal surfaces exposed to an external electric field become charged. In metals, it is known that the electrostatic field does not penetrate deeply beneath the surface (only within 1-2 monoatomic layers [63]) because free electrons re-distribute within the metal bulk to screen the external electric field at the surface. Hence, it is sufficient to understand the nature of the surface charge in terms of excess or depletion of electron density at the surface atoms depending on the direction of the applied electric field. Generally speaking, the redistribution of electron density is not uniform and depends on the positions of the atoms with respect to the positions of the surrounding surface atoms. This results in a non-uniformity of the partial charges that can be assigned to individual surface atoms. The partial charge associated with the atom responds to the applied electric field, adding the forces of electrostatic nature to the a dynamics that may modify the evolution of the surface morphology under the applied field.
[…]
@@ V.2 Work function variation due to surface irregularities
where is the current density and is the work function. is the local electric field that is estimated as , where is the applied electric field and the enhancement factor which is obtained from the fits to the experimental data in order of 20-80 [76]. It is apparent from this formula that it is often convenient to plot experimental data as as a function of , as is illustrated for example in Fig. 30.
[…]
The minor reduction of the work function due to these geometric asperities on the surface was assessed to add a factor of 1.14 and 1.24 [168] to the apparent field enhancement in case of a single adatom and a pyramid as compared to the flat Cu surface, respectively. This illustrates that the geometric protrusion can cause the simultaneous reduction of the work function, which can be seen as a possible factor contribution to the large values of the field enhancement factors measured in experiment [76]. A similar conclusion was derived in the experimental study [25], where a synergistic effect between the reduction of work and a reasonable geometric field enhancement could explain the measured high values of the field emission currents. However, it is clear that the possible reduction of the work function alone cannot explain the high values of field enhancement factors measured experimentally, still necessitating the assumption of a purely geometric field enhancement to explain the experiment.
@@ V.3 Role of electron density distribution on surface evolution under external electric fields
To explore the long-term mechanisms of growth of surface asperities, it is necessary to include in the modeling of surface evolution under external electric fields the forces that the field exerts on the charged metal surface.
@@
First, the interaction of the charged surface with the field gives rise to Maxwell stress within the material, which is always tensile independently of the direction of the field [63]. Because the charge distribution is not uniform at surface asperities, the stress in these regions is also not uniform. Furthermore, the continuum Maxwell stress at the atomic scale can be seen as a Lorentz force acting on the partial charge assigned to individual surface atoms. These forces follow the dynamic changes in the electric field distribution acting stronger on the atoms with higher partial charge. This process alone can trigger dramatic changes on the surface, in particular, when an under-surface defect is present, such as a void or a cavity [140]. Moreover, these forces contribute to field-assisted evaporation of atoms from sharp surface features at elevated temperatures [36].
[…]
A metal surface is known to develop a permanent dipole moment, see, e.g. [84], which is not uniform and may vary around the surface atoms with different numbers of neighbors. This means that the electron densities on metal surfaces are slightly shifted with respect to the position of the positive nuclei of surface atoms. As was discussed earlier (see Section V.1), under an external electric field, the metal surface becomes charged, i.e. the electron densities already shifted (with respect to the nuclei) due to the permanent dipole moment, are shifted either further towards the vacuum or inwards the surface depending on the direction of the applied field. This causes excess or depletion of electron density above the atoms facing the vacuum, which was observed in a number of DFT studies of metal surfaces subject to external electric fields [124, 37, 119].
@@
Field enhancements near surface asperities imply the appearance of field gradients which can be used for modelling of the migration of partially charged atoms on the surface. Since the increase of the electric field increases the partial charge on the surface atom, it can be expected that such partial charge associated with the atom will promote migration of this atom towards stronger fields. The first idea of associating the behavior of partially charged surface atoms with the isolated dipoles was proposed in Ref. [169], where the adatoms were suggested to migrate preferentially toward the surface electric field gradients, similarly to the way isolated dipoles move in the presence of field gradients. This analogy was proposed to explain experiments in which the biased diffusion toward sharp features or edges of sharp tips was observed [171, 72]. Better physical models to explain the biased diffusion in the presence of electric field gradients were based on density functional theory calculations [47]. However, these calculations are too slow to assess the dynamic effect of the biased diffusion under applied electric fields. To enable more systematic observation with statistical analysis of the effect, it is necessary to quantify the effect to enable its incorporation into large-scale models or analytical theories.
[…]
Figure 49a also shows that the adatom induces a significant charge redistribution even in the absence of an external electric field. This introduces an additional dipole moment associated with the adatom. This is a well-known polarization effect typically associated with a local work function change [124]. Moreover, the dipole moment associated with the adatom is dynamic and changes depending on the local atomic environment and the position of the adatom with respect to the lattice sites. Moreover, the dipole moment itself as well as the rate of its change depends on the strength of the external electric field. The interaction of this dipole with the external electric field introduces an electrostatic energy into the system. Depending on how changes along the path of migration of the adatom, the barrier can increase or decrease, as seen in Figure 48.
Moreover, a field gradient introduces additional terms in the migration energies, resulting in the following form for the migration barrier, as derived in Ref [84]
@@
In the above equation, is the migration barrier without a field, and refer to the difference of the effective dipole moment and the effective polarizability between the saddle point (the highest energy point between the lattice sites) of a migration process and the lattice site itself, and is the applied field.
An example of the functional form given by Equation (4) presented in Figure 48 shows the monotonous decrease of the migration barrier with increase of the negative electric field.

### v2

@@ IV.5 Pulse length dependence
In most applications, the electromagnetic field driving the breakdown is a pulsed rf signal. Thus, the driving forces act for a limited pulse length, , which allows for exploration of the kinetics of the breakdown nucleation driving force. Assuming that the plastic activity within the sample follows stochastic evolution leads to a strong dependence of the on , which can be empirically shown to satisfy
@@
see Fig. 45 [48]. Here, is the breakdown rate and , , , and are constants depending on the field. Such behavior was measured in connection with the CLIC project where pulse lengths between 50 and 400 ns were examined, with a duty cycle of 50 Hz [194]. Measurements of the dependence of breakdown rate on pulse length are described in subsection III.5.
@@
In various scenarios, the BDR was shown experimentally, to have an exponential or power-law dependence on [42, 61, 34]. However, the validity of using the existing data to determine the dependence is limited, as it consists of either a small sample [42] or of measurements taken during the conditioning process, when the BDR is still dominated by extrinsic processes [34]. At the very least, the BDR is expected to saturate for a continuous-wave rf signal. Therefore, the exponential or power-law dependence holds only for a limited range of pulse lengths.
@@
Under the assumption of inter-pulse independence, as discussed in section III with relevant data shown in Fig. 17, the distribution of breakdowns in time within each pulse should be an increasing function due to the finite evolution of the time to transition . Measurements of the distribution of time of breakdowns within a pulse are described in Sec. III.3. Figure 46 shows this distribution for = 250 MV/m and a pulse duration of 400 ns. For a given time interval within a pulse, an event will mature if it began within the time interval . Given that a breakdown occurred, the probability that it occurred within an interval is, therefore, , where is linearly proportional to the probability distribution function. This integral, however, is simply the CDF of , which is derived from the MDDF model. This non-Poissonian distribution becomes predominantly Poissonian for times that are significantly greater than .
@@
If, however, the interval between pulses is smaller than the typical relaxation time, then the breakdown probability should not depend on the pulse duration alone, but rather on the combined effect of exposure to the field and the relaxation achieved between pulses. In this case, the variation in the BDR within the pulse can be small and characterized by a constant probability, similarly to the slow variation observed for ns in Fig. 46. Indeed, in [195] it was shown that the breakdown distribution does not vary significantly within the pulse. However, an increase in the breakdown probability was observed for one of the structures in [195], and it was not clear whether full conditioning was achieved. Therefore, experiments involving variation of the duty cycle, as well as additional data regarding the pulse-length dependence during and after conditioning, can help determine the exact nature of the pulse-length dependence of the BDR and help determine the memory effect between and within pulses.
@@
The success of dislocation-based models in describing distributions of BD, and pre-BD phenomena, together with the consistent microscopic observations, provides significant support to the hypothesis that indeed dislocation-based dynamics control the early stages prior to plasma onset.
@@ IV.6 Other indications of plastic activity
Significant dislocation activity can lead to various signals, including electric signals, due to effects on resistivity and acoustic emission due to the collision of dislocations with obstacles. In these cases, macroscopic signals help quantify variations in the global density of dislocations or significant increases in the number of moving dislocations. Such a measurement, independent of microscopic observations, may help corroborate the assumption that dislocation motion is related to the observed processes. For plastically induced cyclic loading, it was shown previously that moving dislocations produce a measurable acoustic emission signal [104]. However, efforts to measure the acoustic emission for surfaces exposed to electric fields have not yet produced measurable signals, due to the inaccessibility of the emitting area to direct acoustic monitoring. Additional examples of such measurements include the effects of dislocation density on the local variation of resistivity in the top layer, which is expected to affect the resistivity of the uppermost grains and yield a measurable signal, as observed in direct plastic deformation [126]. For both methods, initial trials were performed, and for the resistivity system, a detailed design was simulated [28], but as of now, the measurability of pre-BD signals remains an unsolved issue.
@@ V Surface dynamics
It is customary to associate the vacuum breakdown activity with the various surface irregularities of different origin that may give rise to the field emission currents eventually developing into full vacuum arcing events [74, 73, 15, 6, 98, 99]. Mainly these irregularities are associated with surface contamination or post-arcing sharp feature at rims of cathode spots. Alternative hypothesis of a potential whisker growth to serve as an initial field emitting tip, has also been proposed [101]. To date, however, there is no consistent theory which can show the relevance of these processes to the pre-breakdown condition on metal surfaces exposed to high electric fields. To understand better the physics of metal surface evolution in the presence of a high electric field, it is important to take a deeper look into materials properties and the mechanisms that may govern modification of these properties under applied electric fields.
@@
Surfaces subject to high electric fields undergo several dynamic processes that can create and/or amplify various surface irregularities. As was shown in Section IV, some few atomic layer protrusions, step edges, or even larger irregularities may develop on the surface as a result of under-the-surface dislocation dynamics triggered by the stress that is a consequence of the application of an external electric field.
@@
Strictly speaking, the interatomic forces that hold the material together at the surface act to smooth out any irregularities on the surface following the principle of minimization of surface energy [24]. This implies that any atomically sharp surface feature experiences a strong thermodynamic force towards a blunter and, hence, more energetically stable shape. This effect was observed in kinetic Monte Carlo atomistic simulations in Ref. [68]. However, the features resulting from the dislocation activity under the surface can develop further as a result of electrostatic forces acting on partially charged surface atoms by the applied electric field. Although the electric field itself does not penetrate below the first atomic layer, these forces translate into tensile stress throughout the metal bulk because of the interatomic interactions. Because of the nature of these interactions, the force on surface atoms is always directed out from the surface, imposing tensile stresses within the material. These stresses can trigger dislocation dynamics that leads to the growth of asperities on the surface of the metal exposed to a high electric field [149].
@@
There is significant experimental evidence that in the presence of an external electric field, the diffusion of surface atoms becomes biased towards field gradients (i.e. inhomogeneity of the electric field around surface irregularities) [180, 179, 77]. This effect was shown to lead to the formation of surface features; for example, nanoprotrusions were seen to grow in several experimental studies [17, 131, 57, 197, 117].
@@
Figure 47a illustrates the faceted surface of a W tip exposed for 5 hours to the simultaneous effect of an electric field and local heating by a laser [197]. The inset shows the field emission map after 2.5 hours from the beginning of the experiment. Here, the bright spots correspond to the localized field emission currents as a result of the appearance of field-emitting spots along the edges of the surface facets. Weak emission is still visible from the ridges, which was stronger at the beginning of the experiment but started fading away when more localized field-emitting spots started to grow. Eventually, several initial spots merged into a single one at the sharpest corner (similar to that shown in Fig. 47b) with a much stronger current at a lower applied voltage. We note that the W tip was carefully relaxed prior to the experiment to ensure the absence of any irregularity on the tip surface, which could trigger the development of the observed faceted structures with the subsequent formation of field-emitting spots.
The significant reduction of the applied voltage (up to three times) observed in this experiment, which still led to a strong field emission signal, cannot be explained by any other phenomenon. The corresponding finite-element method (FEM) calculations, see Fig. 47b, confirmed that similar field enhancement (a.k.a. the applied voltage reduction) can be obtained if one assumes the formation of a protrusion at the corner of a surface facet. The latter may form as a result of thermal stresses during rapid heating and cooling of the surface by femtosecond laser irradiation. Here, the color bar indicates the electrostatic field value with the highest value changing from left to right from 2 V/nm to 7.34 V/nm, measured at the small corner protrusion.
@@ V.1 Theoretical modeling of surface charge effects on atomic dynamics
Metal surfaces exposed to an external electrostatic field become charged but the field does not penetrate deeply beneath the surface, only within 1-2 monoatomic layers [66], because free electrons re-distribute within the metal bulk to screen the external electric field at the surface. Hence, it is sufficient to understand the nature of the surface charge in terms of excess or depletion of electron density at the surface atoms depending on the direction of the applied electric field. Generally speaking, the redistribution of electron density is not uniform and depends on the positions of the atoms with respect to the positions of the surrounding surface atoms. This results in a non-uniformity of the partial charges that can be assigned to individual surface atoms. The partial charge associated with an atom responds to the applied electric field, adding electrostatic forces to the atomic dynamics that may modify the evolution of the surface morphology under the applied field.
The first approach to tackle this problem was developed by [39]. They proposed an algorithm for assigning partial charges (i.e. a fraction of an electron) using the Gaussian pillbox to estimate the charge induced by the local field according to Gauss’s law at every atom [66]. In this way, each individual surface atom is assigned a partial charge depending on its position at the surface and the local atomic environment. In this work, the distribution of the electrostatic field in vacuum is calculated by solving the Laplace equation for the applied field imposed as a boundary condition far above the surface. The solution is updated on the fly following the dynamic change in the surface morphology. A snapshot series in Fig. 48(a) illustrates atomistic simulations of the resulting dynamic evolution of a nanometer-size tip on the Cu surface under an electric field.
Slow diffusion processes are very challenging for molecular dynamics simulations, which can only cover very short time spans of a few nanoseconds [171]. To allow any dynamics to be observable within this short time span, elevated temperatures (600 K) and electric fields ( 8-10 GV/m) are needed. The atoms are colored according to the partial charge induced on the surface atoms by the applied electric field. The latter is enhanced around the tip as can be seen in Fig. 48(b), where the cross-sectional views of the same surface tip in Fig. 48(a), including the equipotential contour lines to indicate the field enhancement evolution following the shape modifications of the surface feature. As one can see, the partial charge in these simulations is sufficient for field-assisted evaporation of the top atoms already within the short time span of the molecular dynamics simulations. The dependence of the partial charge value of a surface atom on its immediate environment has been verified by employing density functional theory (DFT) calculations for a single and a couple of adjacent Cu adatoms on a Cu surface [40]. In the context of the theoretical model, the adatom is an atom that lies above the top layer of a crystal and has fewer atomic bonds to the surface atoms than the surface atoms themselves within a complete layer. As such, adatoms are exposed to vacuum and, hence, to an applied electrostatic field stronger than regular surface atoms.
a)
b)
@@ V.2 Work function variation due to surface irregularities
where is the current density and is the work function. is the local electric field that is estimated as , where is the applied electric field, and is the enhancement factor obtained from the fit to the experimental data in order of 20-80 [80]. It is apparent from this formula that it is often convenient to plot the experimental data as as a function of , as illustrated, for example, in Fig. 31.
Since high sharp features have not yet been reported on regular surfaces, there is also a suggestion that a strongly increased field emission current density may be caused by local variations in the work function ( eV [8]) due to surface roughness and various surface defects [169]. This is a plausible assumption since, according to the Fowler-Nordheim theory, the reduction of the work function causes the same effect as the enhancement of the local electric field, with the two being practically indistinguishable in a current-voltage measurement. Moreover, is well known to be sensitive to the crystal orientation of the surface [167], as well as to defects. The difference between ridged and closed surfaces in FCC structures may reach a few tenths of eV [49, 18]. Even the vicinal surface can change the work function by a couple of percent [59]. Moreover, photoemission spectroscopy reported a local drop of nearly 0.1 eV of the work function at the step edges on atomically smooth surfaces relative to the measured overall value [169]. However, strong plastic deformations were suggested to also alter the value of the work function [105, 107].
In computational modeling, this effect can be accurately assessed using density functional theory calculations, where the electron density distribution around the surface atoms is calculated explicitly. Because of the limited length scale in these calculations, the analysis of only small defective structures on a metal surface is feasible. However, some information about the variation of the work function around surface defects can be deduced from these calculations. For instance, despite insignificant dimensions, a single adatom on a flat surface can already contribute to the surface roughness, causing a local disturbance to the electric field distribution around it. This disturbance causes a local decrease in the dipole layer, which in turn is expected to modify the work function. Earlier DFT calculations showed that a drop in work functions can be expected due to the presence of a single adatom on the copper surface , and it was found to be as large as 5.6% in the very close vicinity of the defect that smeared at 1. 5% while averaging the potential energy curve on a larger surface area [40]. Later, similar calculations, but for the Cu surface [178], confirmed a similar reduction of the work function in the presence of an adatom, a monoatomic step. A further reduction of up to 10% was found assuming the growth of a pyramid on the surface. However, the calculations suggested that no further decrease in the work function can be expected since the dominant reason for this decrease is the maximal redistribution of electron densities around the top surface atom in a position which is the most exposed to the vacuum. Any other position of the atoms on the surface will result in less exposure (i.e., more neighboring atoms on the surface) and hence in lower effect on the work function.
The presence of geometric asperities on the surface was determined to induce a minor decrease in work function, contributing multiplicative factors of 1.14 and 1.24[178] to the apparent field enhancement in the case of a single adatom and a pyramid. This illustrates that geometric protrusion can cause simultaneous reduction of the work function, affecting field enhancement factors measured in experiments [80]. A similar conclusion was derived in the experimental study [27], where a synergistic effect between reduction in the work function and reasonable geometric field enhancement could explain the high values of measured field emission currents. However, it is clear that the possible reduction of the work function alone cannot explain the high values of field enhancement factors measured experimentally, still necessitating the assumption of a purely geometric field enhancement to explain the experiment.
@@ V.3 Role of electron density distribution on surface evolution under external electric fields
Surface asperities, even single adatoms, i.e. the atoms that randomly appeared on the top of the surface, modify the electron density landscape on the surface. The distribution of non-uniform electron density due to the presence of an adatom on the tungsten (W) surface can be seen in Fig. 49a. Here, the results of quantum-mechanical density functional theory (DFT) calculations [89] are shown for a W atomic structure with a W adatom in the middle of it, viewed from the top. The position of the adatom with respect to the surface atoms is seen in Fig. 49b, where the side view of the same geometry is shown. Although the nature of the chemical bonds is the same for all atoms in this system, the smaller number of neighbors of the adatom reduces its binding to the surface. The position of the adatom also promotes accumulation of the electron density right above it (cyan cloud in the middle). This electron density redistribution occurs already in the absence of an external electric field. The excess of electron density around the adatom creates a localized alternation of the electron density around the surface atoms beneath the adatom, which can be seen in Fig. 49a as magenta and cyan colors that, respectively, show depletion and excess of electron density . As discussed in Section V.2, this effect is responsible for a minor local reduction in the work function [133, 178].
Figure 49b shows that under a positive external electric field (anode), the surface is positively charged. The color around the surface atoms is now magenta, while the bottom of the atomic structure used in the DFT calculations is colored in cyan, indicating the negative charge of this surface. (Note that in static calculations no electric current is included.) Again we see that the adatom is responsible for the non-uniform charge distribution. Although in this case, all surface atoms become charged, the charge associated with the adatom is much higher. This non-uniformity is expected to play an important role in the evolution of the surface morphology for metals exposed to high electric fields, since the interaction of the atoms with the electric field will be different in different surface locations.
@@
First, since the interaction of the charged surface with the field generates tensile stresses also under the surface [39], the nonuniform charge around the surface asperities will give rise to the stress concentration underneath these regions where the charge is higher. These forces follow the dynamic changes in the electric field distribution that act more strongly on the atoms with higher partial charge. This process alone can trigger dramatic changes on the surface, in particular, when an under-surface defect is present, such as a void or cavity [149]. Moreover, these forces contribute to field-assisted evaporation of atoms from sharp surface features at elevated temperatures [39].
On the other hand, the non-uniform charge distribution on surface atoms may lead to the evolution of surface morphology, which proceeds via surface diffusion processes. These thermally activated hoppings of surface atoms and adatoms from one lattice site to another [7] will be affected by an instantaneous change of the charge accumulated on the atoms that perform hoppings due to the interaction of the induced partial charge on the surface atoms with the applied electric field. These energy barriers are known as migration barriers that define the hopping rates. To incorporate this effect within computational models, one can also consider the interaction of a charged metal surface with the applied electric field as follows.
A piece of metal is known to develop a macroscopic permanent dipole when exposed to an external field. The permanent dipole moment is defined by the accumulated charge on the metal surfaces that compensates for the electric field within the metal itself. Microscopically, this process can be seen as charge accumulation on surface atoms, which is, in turn, not uniform and depends on the configuration of surface atoms. Depending on the direction of the applied electric field, the atoms facing the vacuum will have excess or depletion of electron density above them [133, 40, 128], and the stronger the atoms stand out from the surface, the higher the partial charge around them.
@@
Field enhancements around surface asperities imply appearance of field gradients which can be used for modelling the migration of partially charged atoms on the surface. Because the increase of the electric field increases the partial charge on the surface atom, it can be expected that such a partial charge associated with the atom will promote migration of this atom towards stronger fields. The first idea of associating the behavior of partially charged surface atoms with the surface dipoles was proposed in Ref. [179], where the adatoms were suggested to migrate preferentially toward the surface electric field gradients, similarly to the way isolated dipoles move in the presence of field gradients. This analogy was proposed to explain experiments in which the biased diffusion toward sharp features or edges of sharp tips was observed [181, 76]. Better physical models to explain biased diffusion in the presence of electric field gradients were based on DFT calculations [50]. However, these calculations are too slow to assess the dynamic effect of the biased diffusion under applied electric fields. To enable more systematic observation with statistical analysis of the effect, it is necessary to quantify the effect to enable its incorporation into large-scale models or analytical theories.
Recently, a consistent theory of the polarization of a metal surface and its localized effect on surface atoms and adatoms was proposed in Ref. [89]. In this work, the polarization of the macroscopic surface is considered as a whole, while significant charge redistribution around a surface feature (an adatom in Fig. 49a) is taken into account as an additional dipole moment associated with it. This additional dipole moment exists even in the absence of an external electric field. It has also been seen that the additional dipole moment depends on the local atomic environment and the position of the atom with respect to the lattice sites.
The dipole moment itself and the rate with which it can change depend on the strength of the external electric field. The interaction of this dipole with the external electric field introduces an additional energy component into the system, which is an electrostatic energy . Figure 50 demonstrates this effect and how the energy modifies the binding energies (green) and migration barriers (black) as a function of the field strength. These DFT calculations clearly show the reduction in the binding energy to the surface (right-hand -axis) and the corresponding decrease in the migration barrier (left-hand -axis) for negative fields. However, some increase in these values for positive electric fields is associated with the compensation of the additional permanent dipole moment generated at the adatom in the absence of the field. At low positive fields, the charge is pressed inwards, reducing the dipole moment and, hence, decreasing the effect of the electric field on the height of the barrier.
Moreover, this consideration can take into account the effect of the field gradient, which can be introduced as an additional term in the migration barrier calculations as follows [89]:
@@
In Eq. (6) is the migration barrier without a field, and refer to the difference in the effective dipole moment and the effective polarizability between the saddle point (the highest energy point between the lattice sites) of a migration process and the lattice site itself, and is the applied field. Having these considerations in mind, one can attempt to simulate surface diffusion processes, explicitly including the effects of electrostatic fields in the computational models.